The sum of the two roots of a quadratic equation is 5 and their product is -6. What could be the equation?
Answer by ankor@dixie-net.com(22740)

The sum of the two roots of a quadratic equation is 5 and their product is -6. What could be the equation?
:
let the two roots be a and b
a + b = 5
b = (5-a)
and
ab = -6
so we have
a(5-a) = -6
5a - a^2 = -6
0 = a^2 - 5a - 6 could be the equation
Factors to
(a-6)(a+1)= 0
two solutions
a=6, then b=-1
a=-1 then b=6